Find an equation of a line passing through the points (8,10) and (15,8)
Please help I've tried to solve it and I can't find the right answer.

Answer by stanbon(75887)

Find an equation of a line passing through the points (8,10) and (15,8)
-------
Form: y = mx + b
----
Find "m" and "b"::
----
m = (8-10)/(15-8) = -2/7
----
Substitute for y,x, and m::: solve for "b":
10 = (-2/7)8 + b
(70/7) = (-16/7) + b
b = 86/7
-----------------
Equation:
y = (-2/7)x + 86/7
====================
Cheers,
Stan H.